OpenStudy (anonymous):

Which fraction is equivalent to 5%? A.1/2 B.1/20 C.50/100 D.10/50

OpenStudy (kendricklamar2014):

The answer is actually B

OpenStudy (kendricklamar2014):

because 1/20 = 0.05 * 100 = 5%

OpenStudy (anonymous):

yes its b because 5% = 0.05 my bad

OpenStudy (welshfella):